OpenStudy (anonymous):

5(8 - n) = 3n-16 Please show me the steps on how to solve thanks

OpenStudy (anonymous):

distribute and collect like terms.

OpenStudy (anonymous):

could you show me step by step? so i can see it, i dont know how.

OpenStudy (anonymous):

40-5n=3n-16 -5n-3n=-16-40 -8n=-56 n=7

OpenStudy (anonymous):

Ty